# RETTA: Retrieval-Enhanced Test-Time Adaptation for Zero-Shot Video Captioning

### v2

@@ Abstract
Despite the significant progress of fully-supervised video captioning, zero-shot methods remain much less explored. In this paper, we propose a novel zero-shot video captioning framework named Retrieval-Enhanced Test-Time Adaptation (RETTA), which takes advantage of existing pre-trained large-scale vision and language models to directly generate captions with test-time adaptation. Specifically, we bridge video and text using four key models: a general video-text retrieval model XCLIP, a general image-text matching model CLIP, a text alignment model AnglE, and a text generation model GPT-2, due to their source-code availability. The main challenge is how to enable the text generation model to be sufficiently aware of the content in a given video so as to generate corresponding captions.To address this problem, we propose using learnable tokens as a communication medium among these four frozen models GPT-2, XCLIP, CLIP, and AnglE. Different from the conventional way that trains these tokens with training data, we propose to learn these tokens with soft targets of the inference data under several carefully crafted loss functions, which enable the tokens to absorb video information catered for GPT-2. This procedure can be efficiently done in just a few iterations (we use 16 iterations in the experiments) and does not require ground truth data. Extensive experimental results on three widely used datasets, MSR-VTT, MSVD, and VATEX, show absolute 5.1$\sim$32.4improvements in terms of the main metric CIDEr compared to several state-of-the-art zero-shot video captioning methods.
@@
[…]
@@ 1 Introduction
Video captioning, aiming to automatically describe the video content in one sentence, is one of the traditional vision-language tasks. It plays a crucial role in numerous applications, such as video title generation [1], blind assistance [2], and video search [3]. Although current video captioning research makes remarkable progress on full-supervision algorithms, exploration in the area of zero-shot learning still remains limited.
Recently, with the development of pre-trained foundation models in vision and language areas, their combination has shown outstanding zero-shot learning capabilities. For example, for the text-based question-answering task, pre-trained retrieval models have been integrated with Pre-trained Large language Models (PLMs) to improve the quality of generated answers. An example is shown in Fig. 1 (a), where the retriever takes the question “The 2024 Olympics will be held” as a query to retrieve relevant knowledge “The 2024 Summer Olympics, known as Paris 2024, will be hosted in Paris, France, from 26 July to 11 August.”. By concatenating it with the question as a prompt for the model, we get the output “Paris, France, from 26 July to 11 August 2024”.
@@
However, a simple application of this paradigm does not work well for multi-modal visual description tasks (e.g., video captioning). For example, as shown in Fig. 1 (b), the video query retrieves the related sentence “Cutting a piece of bread”; then together with the hard prompt “Video showing” serving as the input of PLM, the PLM outputs “how to cut”, which does not meet the expectation of video captioning. The main reasons lie in two aspects: (1) Weak semantic connection: The semantic connection between the prompt “video showing” and the retrieved text is not as strong as that between questions and answers in QA tasks. This hampers PLMs to generate expected outputs. (2) Improper prompts: Prompt engineering itself is a challenging task, letting alone to additionally take diverse retrieved text into consideration.
@@
To address the above-mentioned problems, we propose to leverage learnable tokens as a connection medium among different foundation models. These tokens should convey video information to PLMs and indicate to PLMs what type of output is expected. Moreover, we expect the combination of these pre-trained models is able to quickly adapt to the task, preferably without the need of training. To meet the quick adaptation requirement, we propose to update these tokens during inference but with soft targets, which gets rid of the conventional training of soft prompts. Our experiments show that the whole framework can well adapt to the video captioning task after just a few of iterations (we use 16 iterations). To meet the former content awareness and output type requirements, we design several loss functions aligning the generated text and soft targets. Specifically, the soft targets are calculated by retrieved text and candidate sentences to guide the choice of candidate words at each time step. This is because soft targets can help the model consider multiple potential tokens in text generation, boosting the diversity of the generated results and improving generalization. An example is shown in Fig. 1(c), where our method leads to better video captions.
@@
  • Extensive experimental results on three widely-used datasets show that our method outperforms several state-of-the-art zero-shot video captioning approaches with absolute 5.1$\sim$32.4improvement in terms of the main metric CIDEr.
@@ 2 Related work
Video Captioning Recently, a series of full-supervision video captioning research [4, 5, 6, 7, 8, 9, 10, 11, 12] have made significant progress. However, fewer efforts have been made on the video captioning task under the zero-shot setting [13, 14, 15, 16]. ZeroCap [13] uses a visual-language model to guide a language model to describe the given image via the designed CLIP-loss. EPT [14] further extends it to video captioning using the proposed CLIP-loss to optimize the pseudo-tokens. To improve inference speed, MAGIC [15] optimizes the decoding strategy of language models and employs CLIP scores to adjust PLM logits for better alignment with image correspondence. DeCap [17] and MultiCapCLIP [18] also leverage the visual-language alignment ability of CLIP for zero-shot video captioning. They first utilize text-only data to train a decoder from scratch and then input video to generate captions at the inference stage. Although the above-mentioned methods improve zero-shot caption quality, they still face the challenge of understanding video content. To tackle this problem, we propose a retrieval-augmented video captioning framework that retrieves video-related knowledge to deepen the understanding of the video.
Video-Text Retrieval Video-text retrieval (VTR) aims to find the most relevant text based on a video query. To retrieve the correct video candidates given a text query, recent works [19, 20, 21, 22, 23, 24] focus on the cross-modal alignment between the related videos and text samples. Most of them can be classified into coarse-grained alignment or fine-grained alignment. The coarse-grained alignment methods mainly leverage the frame-level or video-level visual features to perform alignment between the video and query features. Clip4clip [19] transfers the image-text knowledge of the CLIP model to the VTR task in a coarse manner. X-Pool [20] designs a cross-modal attention model that performs coarse-grained alignment between a text and the frames of a video. TS2-Net [21] introduces the token shift and selection transformer to better model video representation in both temporal and spatial dimensions. However, the coarse-grained alignment is not suitable for capturing detailed correspondence between texts and videos. To overcome this challenge, TokenFlow [22] presents a fine-grained alignment for token-wise similarity calculation. X-CLIP [23] adopts multi-grained contrastive learning to boost video-text retrieval performance, covering fine-grained, coarse-grained, and cross-grained contrast. Another X-CLIP [24] extends CLIP for general video-language understanding using the Kinetics-600 [25] dataset. In this work, we adopt [24] as our retriever as its training data has no overlap with video captioning datasets.
@@ 3 Method
As shown in Fig. 2, our framework contains four pre-trained frozen models: XCLIP [24] and CLIP [31] provide video-level and frame-level video understanding; GPT-2 [32] provide the text generation capability; AnglE [33] is used to calculate text-text similarity. According to their functions for the video captioning task, these pre-trained models can be grouped into a video-text retrieval module and an augmented caption generation module. Specifically, XCLIP serves as the video-text retrieval module, providing a basic understanding of the video. It retrieves video content-related sentences from training text corpus, and we highlight its high-frequency words, as detailed in Section 3.1. The remaining pre-trained models compose the augment caption generation module, which utilizes several test-time learnable tokens to connect all these pre-trained models, absorb video information, and serve as proper prompts for text generation as detailed in Section 3.2. In Section 3.3, we provide more details of the inference process.
@@ 3.1 Video-Text Retrieval
The video-text retrieval module adopts a multi-modal retriever XCLIP to retrieve video-related descriptions from a text corpus. XCLIP is pre-trained on the Kinetics-600 [25] dataset and has the zero-shot ability for video classification and video-text retrieval tasks. Traditional retrievers consistently fetch the top-$k$ ranked sentences. However, this retrieval strategy faces two challenges for video captioning. First, choosing the hyper-parameter $k$ is challenging since the optimal $k$ varies across different videos. Second, retrieving a fixed number of $k$ sentences may introduce noise, as not all of these sentences are closely related to the video.
@@ 3.2 Augmented Caption Generation
As shown in Fig. 2, all model parameters are frozen, and only the learnable soft prompts are optimized during the inference process with soft targets to guide GPT-2 to generate text in line with the content of the given video. Take the $i-{th}$ step as an example, the text generation can be described as:
@@
To utilize the retrieved sentences, we transfer them into the above auto-regression process with our proposed retrieval-based loss $\mathcal{L}_{retrieval}$. It consists of two granularities: the sentence-granularity $\mathcal{L}_{retrieval}^{S}$ and the word-granularity $\mathcal{L}_{retrieval}^{W}$ Specifically, the $\mathcal{L}_{retrieval}^{S}$ is formulated as
@@ 3.3 Inference
The caption $s^{*}$ with the highest similarity is selected as the reliable caption. However, due to the limitations of zero-shot learning, the selected caption still contains some noise and lacks fluency. We design a simple but effective sentence-cleaning strategy, which consists of two steps: (1) extracting essential elements like verbs and nouns from the selected sentence, and (2) employing a pre-trained keyword-sentence model CommonGen [34] to reconstruct these elements into a well-organized sentence. CommonGen is a state-of-the-art model for constrained text generation, which takes keywords (such as “knife”, “chopping”, “onion” and “kitchen”) as input and generates a more fluent sentence “a man is chopping onions with a knife in a kitchen”, as shown in Fig. 3. In the end, the cleaned sentence is used as the final result.
@@ 4.1 Datasets and Metrics
Metrics. In our experiments, we adopt four commonly used metrics to evaluate caption quality, including BLEU@4 (B@4) [38], METEOR (M) [39], ROUGE-L (R) [40], and CIDEr (C) [41]. Among these metrics, CIDEr is the main metric.
@@ 4.2 Implementation Details
We use the pre-trained XCLIP [^1] https://huggingface.co/microsoft/xclip-base-patch16-kinetics-600-16-frames as our video retriever. We uniformly sample 16 frames per video for video retrieval. The visual feature extraction and text-image similarity calculation are performed using the pre-trained CLIP ViT-L/14222https://github.com/openai/CLIP model. We first sample three frames every second for the given video. To avoid using duplicate frames, we further use the CLIP-based strategy to subsample the keyframes. We start with the first frame as an anchor. The next frame is included in the set $\mathcal{F}$ if its dot product with the anchor, calculated by the CLIP image encoder, exceeds a specific threshold $\lambda_{CLIP}$. The newly sampled frame becomes the new anchor and the sampling process repeats in the same manner. The $\lambda_{CLIP}$ is set to 0.9 in all experiments. We use the pre-trained GPT-2 medium [^3] https://huggingface.co/openai-community/gpt2-medium as our caption generator, and the pre-trained AngLE [^4] https://github.com/SeanLee97/AnglE to calculate the similarity between the generated caption and retrieval sentences. Last, we leverage the pre-trained CommonGen [^5] https://huggingface.co/mrm8488/t5-base-finetuned-common_gen to clean the generated sentences into fluent ones. The aforementioned pre-trained models are publicly and readily available.
@@
Empirically, $\lambda_{l}$, $\lambda_{v}$, $\lambda_{s}$ and $\lambda_{w}$ are set to 1.6, 1.0, 0.8, and 0.3 respectively. In our method, we set the number of learnable tokens to 5. We set the number of retrieval sentences to 15 and the frequency size to 5. We adopt the AdamW [42] optimizer with a learning rate of 1e-4 and a weight decay weight of 0.3 to optimize our model. Throughout the experiments, we employ 16 generation iterations per video. To prevent sentences from being overly long and repetitive, we limit the number of tokens generated per sentence to 15. We implement our framework with PyTorch and all experiments are conducted on 1 $\times$ RTX-A6000 GPU.
Table 1: Results on MSVD across various models and supervision levels.
@@ 4.3 Performance Comparison
We compare our method with several state-of-the-art models under different settings on three public benchmark datasets. The results are shown in Table 1, 2 and 3. We note that the performance of supervised algorithms is significantly higher than zero-shot methods, but they require expensive annotation costs. Compared to the other zero-shot methods, our method achieves the best performance in terms of the main metric CIDEr on all the three datasets. As shown in Table 1, our method achieves a CIDEr score of 49.8 on the MSVD benchmark, leading by about an absolute 32.4 over the previous best result. On the more challenging benchmark MSR-VTT, our method also boosts the CIDEr score by about 30.6%. On the VATEX dataset, our method achieves a 23.8 CIDEr Score, gaining 27.3% improvement in terms of the main metric CIDEr on VATEX.

### v3

@@ Abstract
[…]
@@
Zero-shot , Retrieval , Video captioning , Test-Time Adaptation
journal: Elsevier
@@ 1 Introduction
Video captioning, aiming to automatically describe the video content in one sentence, is one of the traditional vision-language tasks. It plays a crucial role in numerous applications, such as video title generation [1], blind assistance [2], and video search [3]. However, current video captioning research [4, 5, 6] typically requires full supervision, with at least 10 reference captions per video, leading to high annotation costs. As a practical solution, zero-shot methods reduce the dependence on costly human annotations [7, 8]. This makes them particularly suitable for scenarios where labeled data is scarce.
Recently, with the development of pre-trained foundation models in vision and language areas, their combination has shown outstanding zero-shot learning capabilities [9, 10]. For example, for the text-based question-answering task, pre-trained retrieval models have been integrated with Pre-trained Large language Models (PLMs) to improve the quality of generated answers. An example is shown in Figure 1 (a), where the retriever takes the question “The 2024 Olympics will be held” as a query to retrieve relevant knowledge “The 2024 Summer Olympics, known as Paris 2024, will be hosted in Paris, France, from 26 July to 11 August.”. By concatenating it with the question as a prompt for the model, we get the output “Paris, France, from 26 July to 11 August 2024”.
@@
However, a simple application of this paradigm does not work well for multi-modal visual description tasks [11, 12] (e.g., video captioning). For example, as shown in Figure 1 (b), the video query retrieves the related sentence “Cutting a piece of bread”; then together with the hard prompt “Video showing” serving as the input of PLM, the PLM outputs “how to cut”, which does not meet the expectation of video captioning. The main reasons lie in two aspects: (1) Weak semantic connection: The semantic connection between the prompt “video showing” and the retrieved text is not as strong as that between questions and answers in QA tasks. This is because traditional RAG methods [13, 14] typically use the question as a query to retrieve relevant textual information as a prefix, which effectively guides the model to generate accurate responses. In contrast, in video captioning, when the video itself is used as a query [15, 16], the retrieved text is often weakly related to generic prompts like “video showing”, thereby hindering PLMs from generating the expected outputs. (2) Improper prompts: Prompt engineering itself is a challenging task [17, 18], letting alone to additionally take diverse retrieved text into consideration.
@@
To address the above-mentioned problems, we propose to leverage learnable tokens as a connection medium among different foundation models. Inspired by test-time prompt tuning approaches [19, 20], the learnable tokens are updated during inference using gradient signals from the task loss. Their role is to deliver video-specific information to the PLMs and specify the expected form of output. Moreover, we expect the combination of these pre-trained models is able to quickly adapt to the task, preferably without the need of training. To meet the quick adaptation requirement, we propose to update these tokens during inference but with soft targets [21, 22], which gets rid of the conventional training of soft prompts. These tokens thus serve as a bridge between different frozen models, making them a practical and effective connection strategy in our setting. Our experiments show that the whole framework can well adapt to the video captioning task after just a few of iterations (we use 16 iterations). To meet the former content awareness and output type requirements, we design several loss functions aligning the generated text and soft targets. Specifically, the soft targets are calculated by retrieved text and candidate sentences to guide the choice of candidate words at each time step. This is because soft targets [21, 22] can help the model consider multiple potential tokens in text generation, boosting the diversity of the generated results and improving generalization. An example is shown in Figure 1(c), where our method leads to better video captions.
@@
[…]
@@ 2 Related work
Video Captioning Recently, a series of full-supervision video captioning research [23, 24] have made significant progress. However, fewer efforts have been made on the video captioning task under the zero-shot setting [25, 26]. ZeroCap [25] uses a visual-language model to guide a language model to describe the given image via the designed CLIP-loss. EPT [26] further extends it to video captioning using the proposed CLIP-loss to optimize the pseudo-tokens. To improve inference speed, MAGIC [27] optimizes the decoding strategy of language models and employs CLIP scores to adjust PLM logits for better alignment with image correspondence. DeCap [8] and MultiCapCLIP [7] also leverage the visual-language alignment ability of CLIP for zero-shot video captioning. They first utilize text-only data to train a decoder from scratch and then input video to generate captions at the inference stage. Although the above-mentioned methods improve zero-shot caption quality, they still face the challenge of understanding video content. In addition, using CLIP in open-world scenarios may lead to degraded performance due to attention misalignment and bag-of-words behavior [28]. These issues introduce biases in multi-object scenes, where CLIP tends to focus on larger objects and overlook text order. This suggests that relying solely on CLIP may be insufficient for tackling open-world tasks such as zero-shot video captioning. To tackle these problems, we propose a retrieval-augmented video captioning framework that retrieves video-related knowledge to deepen the understanding of the video.
Video-Text Retrieval Video-text retrieval (VTR) aims to find the most relevant text based on a video query. To retrieve the correct video candidates given a text query, recent works [16, 29] focus on the cross-modal alignment between the related videos and text samples. To facilitate this alignment, some works leverage pre-trained models such as CLIP as the backbone to obtain rich cross-modal representations. For example, Clip4clip [30] transfers the image-text knowledge of the CLIP model to the VTR task in a coarse manner. To better adapt CLIP to the video domain, X-CLIP [31] extends CLIP for general video-language understanding using the Kinetics-600 dataset. In this work, we adopt [31] as our retriever as its training data has no overlap with video captioning datasets.
@@ 3 Method
As shown in Figure 2, our framework contains four pre-trained frozen models: XCLIP [31] and CLIP [36] provide video-level and frame-level video understanding; GPT-2 [37] provide the text generation capability; AnglE [38] is used to calculate text-text similarity. According to their functions for the video captioning task, these pre-trained models can be grouped into a video-text retrieval module and an augmented caption generation module. The video-text retrieval module retrieves relevant sentences from the corpus and highlights key high-frequency words. These elements serve as auxiliary semantic cues for the caption generation module, enabling it to generate more accurate and contextually appropriate captions. Specifically, XCLIP serves as the video-text retrieval module, providing a basic understanding of the video. It retrieves video content-related sentences from training text corpus, and we highlight its high-frequency words, as detailed in Section 3.1. The remaining pre-trained models compose the augment caption generation module, which utilizes several test-time learnable tokens to connect all these pre-trained models, absorb video information, and serve as proper prompts for text generation as detailed in Section 3.2. In Section 3.3, we provide more details of the inference process.
@@ 3.1 Video-Text Retrieval
The video-text retrieval module adopts a multi-modal retriever XCLIP to retrieve video-related descriptions from a text corpus. XCLIP is pre-trained on the Kinetics-600 dataset and has the zero-shot ability for video classification and video-text retrieval tasks. Traditional retrievers consistently fetch the top-$k$ ranked sentences. However, this retrieval strategy faces two challenges for video captioning. First, choosing the hyper-parameter $k$ is challenging since the optimal $k$ varies across different videos. Second, retrieving a fixed number of $k$ sentences may introduce noise, as not all of these sentences are closely related to the video.
@@ 3.2 Augmented Caption Generation
As shown in Figure 2, all model parameters are frozen, and only the learnable soft prompts are optimized during the inference process with soft targets to guide GPT-2 to generate text in line with the content of the given video. Take the ithi-{th} step as an example, the text generation can be described as:
@@
To utilize the retrieved sentences, we transfer them into the above auto-regression process with our proposed retrieval-based loss $\mathcal{L}_{retrieval}$. It consists of two granularities: the sentence-granularity $\mathcal{L}_{retrieval}^{S}$ and the word-granularity $\mathcal{L}_{retrieval}^{W}$ . Specifically, the $\mathcal{L}_{retrieval}^{S}$ is formulated as
@@ 3.3 Inference
The caption $s^{*}$ with the highest similarity is selected as the reliable caption. However, due to the limitations of zero-shot learning, the selected caption still contains some noise and lacks fluency. We design a simple but effective sentence-cleaning strategy, which consists of two steps: (1) extracting essential elements like verbs and nouns from the selected sentence, and (2) employing a pre-trained keyword-sentence model CommonGen [40] to reconstruct these elements into a well-organized sentence. CommonGen is a state-of-the-art model for constrained text generation, which takes keywords (such as “knife”, “chopping”, “onion” and “kitchen”) as input and generates a more fluent sentence “a man is chopping onions with a knife in a kitchen”, as shown in Figure 3. In the end, the cleaned sentence is used as the final result.
@@ 4.1 Datasets and Metrics
Metrics. In our experiments, we adopt four commonly used metrics to evaluate caption quality, including BLEU@4 (B@4), METEOR (M), ROUGE-L (R), and CIDEr (C). Among these metrics, CIDEr is the main metric.
@@ 4.2 Implementation Details
We use the pre-trained XCLIP [^1] https://huggingface.co/microsoft/xclip-base-patch16-kinetics-600-16-frames as our video retriever. We uniformly sample 16 frames per video for video retrieval. The visual feature extraction and text-image similarity calculation are performed using the pre-trained CLIP ViT-L/14222https://openaipublic.azureedge.net/clip/models/b8cca3fd41ae0c99ba7e8951adf17d267cdb84cd88be6f7c2e0eca1737a03836/ViT-L-14.pt model. Our video frame sampling strategy involves two steps: fixed-rate sampling followed by CLIP-based redundancy filtering. First, we uniformly sample three frames per second from the entire video. To reduce redundancy, we then apply a CLIP-based keyframe selection strategy. Starting with the first frame as an anchor, we compute the dot product similarity between each subsequent frame and the anchor using CLIP image embeddings. If the similarity is smaller than a predefined threshold $\lambda_{CLIP}$ (set to 0.9), the frame is added to the keyframe set $\mathcal{F}$ and becomes the new anchor. This process is repeated for the remaining frames. We use the pre-trained GPT-2 medium [^3] https://huggingface.co/openai-community/gpt2-medium as our caption generator, and the pre-trained AngLE [^4] https://huggingface.co/WhereIsAI/UAE-Large-V1 to calculate the similarity between the generated caption and retrieval sentences. Last, we leverage the pre-trained CommonGen [^5] https://huggingface.co/mrm8488/t5-base-finetuned-common_gen to clean the generated sentences into fluent ones. The aforementioned pre-trained models are publicly and readily available. All the pre-trained models used in our study adopt their original default parameters without any manual modifications, making them directly integrable into our system.
@@
Empirically, $\lambda_{l}$, $\lambda_{v}$, $\lambda_{s}$ and $\lambda_{w}$ are set to 1.6, 1.0, 0.8, and 0.3 respectively. In our method, we set the number of learnable tokens to 5. We set the number of retrieval sentences to 15 and the frequency size to 5. We adopt the AdamW optimizer with a learning rate of 1e-4 and a weight decay weight of 0.3 to optimize our model. Throughout the experiments, we employ 16 generation iterations per video. To prevent sentences from being overly long and repetitive, we limit the number of tokens generated per sentence to 15. We implement our framework with PyTorch 1.13.1 and Transformers 4.32.1. All experiments are conducted on an Intel Xeon Platinum 8375C CPU, 64 GB of RAM, and one NVIDIA RTX A6000 GPU.
Discussion of Model Choice. We follow many existing zero-shot captioning works [10, 9, 8, 25, 7] and adopt CLIP as the visual feature extractor and GPT-2 as the text generator to remain consistent with prior approaches. The CLIP model is pre-trained on 400 million image-text pairs, which plays the vision-language alignment role to map the image and sentence into a unified semantic space. The out-of-box GPT-2 is pre-trained on the large dataset WebText collected from a wide range of text data online. It is composed of several Transformer layers, which can cache the previous key and value tokens to infer the next token autoregressively. Therefore, GPT-2 is capable of generating text based on preceding content. XCLIP is pre-trained on the Kinetics-600 dataset and has the zero-shot ability for video classification and video-text retrieval tasks. AnglE is a state-of-the-art model for calculating semantic textual similarity, which is able to be aware of differences among texts. These pre-trained models offer strong zero-shot performance, are lightweight (parameter range: 194.93M–427.62M), easy to deploy, and well-suited for integration into a unified system for collaborative use. These factors make them a reasonable choice. In our study, X-CLIP is used to retrieve reference sentences based on video content. CLIP is responsible for visual feature extraction, frame sampling, and image-text matching to filter candidate tokens. GPT-2 handles text generation, and AnglE computes the similarity between candidate and reference sentences for token selection.
Discussion of Parameter Settings. In this section, we provide a discussion on several important design choices. For video frame sampling, a fixed sampling rate of 3 FPS is adopted, which balances efficiency and completeness for most videos (typically 15–30 FPS). For high-frame-rate videos (e.g., 60 FPS), a higher rate (e.g., 5 FPS) can optionally be used. As illustrated in Figure 4, setting the CLIP similarity threshold to $\lambda_{CLIP}=0.9$ helps eliminate redundant frames while preserving critical visual content, making it a suitable default choice. The discussion of other hyperparameter choices, such as the number of retrieved sentences, the number of learnable tokens, and the frequency size for the word-granularity loss, will be detailed in the Ablation Study section.
Figure 4: Illustration of our CLIP-based sampling strategy. Selected frames are highlighted in red, while filtered (redundant) frames are marked in green.
Table 1: Results on MSVD across various models and supervision levels. B@4, M, R, and C denote BLEU@4, METEOR, ROUGE-L, and CIDEr, respectively. $\uparrow$ indicates that higher is better. This notation applies to all subsequent tables.
@@ 4.3 Performance Comparison
We compare our method with several state-of-the-art models under different settings on three public benchmark datasets. The results are shown in Table 1, 2 and 3. We note that the performance of supervised algorithms is significantly higher than zero-shot methods, but they require expensive annotation costs. Compared to the other zero-shot methods, our method achieves the best performance in terms of the main metric CIDEr on all the three datasets. As shown in Table 1, our method achieves a CIDEr score of 49.8 on the MSVD benchmark, leading by about an absolute 32.4 over the previous best result. On the more challenging benchmark MSR-VTT, our method also boosts the CIDEr score by about 30.6%. On the VATEX dataset, our method achieves a 23.8 CIDEr Score, gaining 27.3% improvement in terms of the main metric CIDEr on VATEX. While our method underperforms slightly on traditional textual metrics such as BLEU@4, METEOR, and ROUGE-L compared to DeCap-COCO and MultiCapCLIP on MSRVTT and VATEX. This is primarily because the latter methods are trained on the full set of reference captions from the MS COCO dataset, thus benefiting from more extensive language supervision. These metrics are better suited for machine translation and may not fully reflect the relevance of captions to video content. In contrast, our task focuses on describing video content, and the notable improvement in CIDEr—a metric specifically designed to measure vision-text relevance—demonstrates the effectiveness of our method in video captioning. Overall, our method performs best on MSVD (CIDEr 49.8), followed by MSRVTT (24.3) and VATEX (23.8). The performance differences mainly stem from the varying complexity of the datasets: MSVD is relatively simple with 1,970 video clips; MSRVTT includes around 10,000 videos with greater diversity; and VATEX, the most challenging, contains 34,991 videos covering 600 action categories. Our method achieves the strongest CIDEr scores across all three datasets, with particularly notable performance on MSVD (49.8). While MSRVTT and VATEX present greater diversity and complexity, our approach still achieves the best CIDEr scores (24.3 and 23.8, respectively), showing robust generalization. This consistent improvement on CIDEr demonstrates that our method effectively captures video content and generates high-quality captions even under diverse scenarios.
In Table 4, we also provide a quantitative comparison of training compute (estimated by GPU memory consumption during training), parameter size, and inference time between our proposed RETTA and the other zero-shot video captioning methods. Both EPT and our RETTA method are test-time adaptation approaches, unlike DeCap-CC3M, which involves extensive pretraining on textual data ((e.g., 25,197 seconds on CC3M)) but offers very fast inference (0.1s per sample). RETTA skips pretraining and performs adaptation during testing. Despite a longer inference time, RETTA delivers a better CIDEr score (23.8 v.s. 18.4). Thus, DeCap-CC3M is more suitable for scenarios where low inference latency is critical (e.g., real-time applications) and higher training cost is acceptable. Furthermore, although RETTA introduces additional overhead due to the integration of four pre-trained models (with an absolute increase of 5334 MB in training memory, 530.07M in parameter size, and 17.8s in inference time compared to EPT), it achieves a significant improvement in CIDEr score (23.8 v.s. 8.9), demonstrating a favorable trade-off between cost and performance. Therefore, for latency-insensitive scenarios (e.g., offline batch processing), both RETTA and EPT are viable choices. In particular, RETTA is better suited when higher performance is required and a text retrieval corpus is available, such as offline summarization of vertical-domain videos.
Table 4: Efficiency and performance comparison. MB, M, and s represent megabytes of GPU memory usage, millions of model parameters, and seconds of inference time, respectively.
Method GPU Memory Usage (MB) $\downarrow$ Parameters Size (M) $\downarrow$ Training Time (s) $\downarrow$ Inference Time (s) $\downarrow$ CIDEr $\uparrow$
DeCap-CC3M 9656 219.41 25197 0.1 18.4
EPT 4408 782.44 0 42.6 8.9
RETTA(Ours) 9742 1312.51 0 60.4 23.8